What mass of glucose (C6H12O6) is needed to prepare 0.225m C6H12O6 using 250.g g of water?
Answer #1

The molality of a solution is defined as the amount of substance (in mol) of solute dissolved in solution is divided by the mass (in kg) of the solvent.

Molality(M): Number of moles of solute / Weight of solvent in kg

Here, 0.225m = x / 0.250Kg, where x = Number of moles of glucose.

So, glucose = 0.05625 moles.

Molar mass of glucose is 180.15 g/mol

So, the amoung of glucose will be = 0.05625 moles. X 180.15 g/mol - 10.13g
Answer #2

What mass of glucose (C6H12O6) is needed to prepare 0.225m C6H12O6 using 250.g g of water?

Solution:- molality(m) = moles of glucose/mass of water in kg

or, 0.225 = moles of glucose/0.25

or, moles of glucose = 0.05625

Now, molar mass of glucose = 180 g/mole

hence mass of glucose required = moles*molar mass = 10.125 g
Answer #3

Calculate the mass of glucose and water required to make 250g of 25% solution of glucose.
Answer #4

Given: Molality of glucose ( C6H12O6 ) = 0.225m

==> Mass of solvent (water) = 250g

==> We know that Molality of glucose = ( Moles of glucose x 1000 ) / Mass of water

==> 0.225 = ( n x 1000 )/250

==> n = ( 0.225 x 250 ) / 1000

==> n = 0.225/4

==> Mass of glucose = No. of moles x Molar Mass of glucose

==> (0.225 x 180)/4

==> 10.125 grams ..........ANS
Answer #5

mass of glucose (C6H12O6) is 4,23 g
Answer #6

435g